Solve the system of equations by graphing. Then classify the system
6x-8y=24
8y-6x=-24
a-the solution is??
b-there are infintely many solutions
c-there is no solution
IS the system consistent or inconsitent?

Answer by mananth(16946)

6x-8y=24
x=0,y=-3 (0,-3)
y=0x=4 (4,0)
y= 3x/4-3
....
8y-6x=-24
x=0,y=-3 (0-3)
y=0,x=-4 (4,0)
y=3x/4-3

b. It is one and the same line. It has infintely many solutions